OpenStudy (anonymous):

What is the hypotenuse of a right triangle that has legs of length 9 units and 13 units?

OpenStudy (mathmale):

Here's where the Pythagorean Theorem comes into play. Please type out that formula. Then we'll use it solve this problem.

OpenStudy (anonymous):

14. But, mathmale is right. Just use the formula!

OpenStudy (anonymous):

a^2 + b^2 = c^2

OpenStudy (mathmale):

Great. Thank you. Now please substitute a=13 and b=9. What do you get? Please type in your result.

OpenStudy (anonymous):

I don't understand how to do it.

OpenStudy (mathmale):

What are the values of \[9^{2}? 13^{2}?\] (That's two separate quantities to calculate.) Do y ou have a calculator handy? Do you know already that 9^2 = 81?

OpenStudy (mathmale):

9 squared = 9^2 = 81. Yes. But 13 squared =13^2 is not 1222. Please try again on that. Have you a calculator?

OpenStudy (anonymous):

81 + 169 = 250

OpenStudy (mathmale):

Great. This is an application of the Pythagorean Theorem. Would you please now find the square root of 250? If done correctly, that will be the length of the hypotenuse of the given right triangle.

OpenStudy (anonymous):

15.8

OpenStudy (mathmale):

that looks fine. Since the 250 in Sqrt(250) can be factored, \[\sqrt{250}=\sqrt{25}*\sqrt{10}=5\sqrt{10}.\] This is equivalent to your 15.8, but in neater form. Feel OK about this problem? if not, what kind of questions have you about it?

OpenStudy (mathmale):

Great working with y ou! Unfortunately, I need to log out of Open Study now; I'm driving from California to Canada, starting this morning. All the best to you. Thanks for the medal. MM